OpenStudy (ny,ny):

The midpoint of AB is M(4,2). If the coordinates of A are (6,-4), what are the coordinates of B?

OpenStudy (anonymous):

The midpoint formula is: \[(\frac{ x _{1} + x _{2}}{ 2 }, \frac{ y_{1} + y_{2}}{ 2 })\] We know that x1 = 6, and y1 = -4 (taken from the coordinates of A) We also know that the midpoint is (4,2) Therefore, \[\frac{ 6 + x }{ 2 } = 4\] and \[\frac{ -4 + y }{ 2}= 2\] Does that make sense?

OpenStudy (ny,ny):

yes.

OpenStudy (ny,ny):

x = 2? and y= 8?

OpenStudy (ny,ny):

so midpoint is (2,8).

OpenStudy (anonymous):

You got it! Glad I could help :)

OpenStudy (ny,ny):

Thanks :)